# Grower-in-the-Loop Interactive Reinforcement Learning for Greenhouse Climate Control

### v2

@@ 1 Introduction
Recently, reinforcement learning (RL) has attracted growing attention in greenhouse climate control. RL is a learning-based method with a model-free nature and adaptive learning capabilities. These features enable RL to perform effectively in complex, error-prone, and uncertain environments [13], which are characteristic of greenhouse systems. As a data-driven approach, RL can continuously optimize control strategies while adapting to dynamic environmental conditions. Several studies have already evaluated RL in simulated greenhouse settings [14, 15, 16], demonstrating its potential for achieving optimal climate control. Despite these promising results, RL applications in greenhouse climate management still face significant challenges. Among the most frequently mentioned limitations are low learning efficiency and limited robustness [16]. Moreover, current RL approaches often exclude growers from the control loop, failing to leverage their domain knowledge and practical experience, which could enhance learning efficiency. This lack of involvement may also lead to growers’ reluctance to adopt RL or other artificial intelligence methods in real-world applications, as they have limited control or understanding of the decision-making process.
@@
    We develop three representative interactive RL algorithms tailored for greenhouse climate control: policy shaping, which uses grower advice on actions; control sharing, and reward shaping, both of which rely on grower’s feedback regarding control performance.
@@
    We analyze how input characteristics are often contradicting, and how the trade-offs between them make grower’s inputs difficult to perfect. We also discuss the potential shortcomings of different methods inputs can be provided.
@@
    We propose a neural network-based approach to enhance the robustness of interactive RL agents under limited input availability by incorporating additional neural networks to aggregate grower’s input.
@@
    We conduct a comprehensive evaluation and comparison of the three interactive RL algorithms in a simulated greenhouse environment, considering the potential variability and characteristics of actual grower input.
@@ 2.1 Greenhouse Environment
where $p_{\text{lettuce}}$ is the price of lettuce (16 Hfl/kg), $\Delta_{\text{dryweight}}$ is the change in lettuce dry weight (kg/m2), $p_{\text{CO}_{2}}$ is the price of CO2 supply (0.42 Hfl/kg), $u_{\text{CO}_{2}}$ is the usage of CO2 supply (kg/m2), $p_{\text{heat}}$ is the price of heating (6.35×10-9 Hfl/J), $u_{\text{heat}}$ is the usage of heating (J/m2), $penalty$ is the cost if constraint violated (5.24×10-3 Hfl/m2). Note that the price of lettuce, heating, and CO2 are obtained from [15]. The penalty is applied when indoor climate constraints [20] (shown in Table 2) are violated. The weather data used in the environment and episode length are further detailed in Section 5.1.
@@
Outdoor CO2 ppm CO2 setpoint ppm CO2 setpoint $-200~{}\mathrm{ppm}$, 0, $+200~{}\mathrm{ppm}$
@@ 2.2 Baseline RL Algorithm
where $\lambda$ is a discount factor. GAE combines multi-step $\delta_{t}$ to balance the trade-off between bias and variance.
@@ 3.3 Simulation of Growers’ Imperfect Inputs
[…]
@@ 4.1 Incorporating Grower’s Inputs
$P(a)=\begin{cases}1-\beta,&a=\pi_{agent}(s,a)\\ \beta,&a=\arg\max_{a}\left(F(s,a)\right)\end{cases}.$ (6)
@@
[…]
@@ 4.2 Addressing Limited Availability
[…]
@@
11:     Update the $F$ function ($\pi_{\text{grower}}$) and $\pi_{\text{error}}$ (if exists) using $D_{\text{input}}$ for $N$ iterations
@@ 4.3 Interactive RL Implementation
[…]
@@ 5.1 Experiment Settings
Learning rate for $\pi$grower network 1×10-4(policy shaping(precise)) / 1×10-3 (reward shaping, control sharing)
Learning rate for $\pi$error network 1×10-3
@@ 5.3 Performance of Interactive RL
Table 9 shows the average cumulative reward of interactive RL algorithms in the test environment. In all experiments, the availability of inputs is full, and the input weights (likelihood) are chosen to be the ones that work best in the test. Additionally, baseline PPO using only simulated feedback as the reward function is also trained and tested, with its test reward reduced to only 0.71 Hfl/m2. This is expected and indicates that the simulated inputs are far from perfect. The policy shaping using precise action advice performs best in this environment, with 8.4% improvement. Policy shaping using constraints and control sharing also increase rewards. However, reward shaping reduces the reward, indicating that it is more sensitive to the quality of input.
@@
Figure 5 shows the average setpoints and indoor climate variables for each hour of the day. All interactive RL algorithms, except for reward shaping, show a similar pattern. They maintain lower temperature setpoints and consequently lower indoor temperatures than the baseline PPO. Notably, indoor temperatures and CO2 setpoints are generally higher than the corresponding indoor values, and high indoor CO2 setpoints and indoor values are observed at night. This is due to several factors: First, the greenhouse environment has no actuator that can actively reduce temperature or CO2 level. Second, in the later stage of the growing cycle, the plant’s respiration can produce a large amount of CO2. Third, since there are no constraints on high CO2 level or temperature at night and relative humidity remains within constraints, the agent is not motivated to use ventilation to lower these variables.
@@ 6 Conclusion
$\phi_{\text{vent,c}}(t)={}$ $\big{(}u_{2}(t)10^{-3}+p_{2,3}\big{)}\big{(}x_{2}(t)-d_{2}(t)\big{)},$
$\phi_{\text{vent,h}}(t)={}$ $\big{(}u_{2}(t)10^{-3}+p_{2,3}\big{)}\big{(}x_{4}(t)-d_{4}(t)\big{)},$
$\phi_{\text{transp,h}}(t)={}$ $p_{4,2}(1-\text{exp}(-p_{1,3}x_{1}(t)))$
×$\times\left(\frac{p_{4,3}}{p_{4,4}(x_{3}(t)+p_{4,5})}\text{exp}\left(\frac{p_{4,6}x_{3}(t)}{x_{3}(t)+p_{4,7}}\right)-x_{4}(t)\right),$
@@
[…]
@@
$x_{4}(t)$ indoor humidity (kg/m3) $d_{4}(t)$ outdoor humidity (kg/m3)
$u_{1}(t)$ CO2 injection (mg/m2/s) $y_{1}(t)$ dry-weight (g/m2)

### v3

@@ 1 Introduction
Recently, reinforcement learning (RL) has attracted growing attention in greenhouse climate control. RL is a learning-based method with a model-free nature and adaptive learning capabilities. These features enable RL to perform effectively in complex, error-prone, and uncertain environments [13] [14], which are characteristic of greenhouse systems. As a data-driven approach, RL can continuously optimize control strategies while adapting to dynamic environmental conditions. Several studies have already evaluated RL in simulated greenhouse settings [15, 16, 17], demonstrating its potential for achieving optimal climate control. Despite these promising results, RL applications in greenhouse climate management still face significant challenges. Among the most frequently mentioned limitations are low learning efficiency and limited robustness [17]. Moreover, current RL approaches often exclude growers from the control loop, failing to leverage their domain knowledge and practical experience, which could enhance learning efficiency. This lack of involvement may also lead to growers’ reluctance to adopt RL or other artificial intelligence methods in real-world applications, as they have limited control or understanding of the decision-making process.
@@
    We developed three representative interactive RL algorithms tailored for greenhouse climate control: policy shaping, which uses grower advice on actions; control sharing, and reward shaping, both of which rely on grower’s feedback regarding control performance.
@@
    We analyzed how input characteristics are often contradicting, and how the trade-offs between them make grower’s inputs difficult to perfect. We also discuss the potential shortcomings of different methods inputs can be provided.
@@
    We proposed a neural network-based approach to enhance the robustness of interactive RL agents under limited input availability by incorporating additional neural networks to aggregate grower’s input.
@@
    We conducted a comprehensive evaluation and comparison of the three interactive RL algorithms in a simulated greenhouse environment, considering the potential variability and characteristics of actual grower input.
@@ 2.1 Greenhouse Environment
where $p_{\text{lettuce}}$ is the price of lettuce (16 Hfl/kg), $\Delta_{\text{dryweight}}$ is the change in lettuce dry weight (kg/m2), $p_{\text{CO}_{2}}$ is the price of CO2 supply (0.42 Hfl/kg), $u_{\text{CO}_{2}}$ is the usage of CO2 supply (kg/m2), $p_{\text{heat}}$ is the price of heating (6.35×10-9 Hfl/J), $u_{\text{heat}}$ is the usage of heating (J/m2), penaltypenalty is the cost if constraint violated (5.24×10-3 Hfl/m2). Note that the price of lettuce, heating, and CO2 are obtained from [16]. The penalty is applied when indoor climate constraints [21] (shown in Table 2) are violated. The weather data used in the environment and episode length are further detailed in Section 5.1.
@@
Outdoor CO2 ppm CO2 setpoint ppm CO2 setpoint $200~\mathrm{ppm}$, 0, $200~\mathrm{ppm}$
@@ 2.2 Baseline RL Algorithm
where GAE trace parameter $\lambda$ combines multi-step $\delta_{t}$ to balance the trade-off between bias and variance.
@@ 3.3 Simulation of Growers’ Imperfect Inputs
Figure 3 illustrates examples of the three input types, focusing solely on temperature. In constraint advice, actions that keep the temperature within the specified range are assigned equal probabilities of selection. Precise action advice sets the temperature as close as possible to the optimal setpoint. Feedback is provided after the outcome of the action ($S\textsubscript{t+1}$) is observed, evaluating whether the resulting temperature is closer to the optimal value compared to the potential effects of alternative actions.
@@ 4.1 Incorporating Grower’s Inputs
[…]
@@
Policy shaping [26] is chosen to incorporate both types of action advices. It influences action selection according to (5). Additionally, a new neural network, $\pi_{\textsubscript{grower}}$, is introduced to estimate the action advice, similar to the $F$ function used in the DQN-TAMER framework.
@@ 4.2 Addressing Limited Availability
In all three algorithms, an additional network ($F$ function or $\pi_{\textsubscript{grower}}$) is used to estimate the input. This network not only facilitates the control sharing method but also enhances the algorithm’s tolerance to limited input availability and high latency. This capability is especially important when inputs must be obtained by querying growers at every step or when providing input demands significant time and effort.
With the additional network ($F$ function or $\pi_{\textsubscript{grower}}$), the goal of providing inputs to the agent’s training is similar to active learning: maximize the learning efficiency of input estimation with minimal input given. As in active learning, two sampling approaches exist: pool-based and stream-based. This paper only focuses on the pool-based approach for two reasons: First, the pool-based approach has better sample efficiency [29], which suits the needs of limited input availability better. Second, the stream-based approach requires immediate decisions (one chance per step), increasing complexity and potentially the burden of the grower if the final decision is given by the grower.
As in active learning, it is essential to efficiently select which steps to provide input. Within pool-based approach, two selection strategies are compared in this paper: a random strategy and a selective strategy. The random strategy samples input steps randomly, while the selective strategy is inspired by SafeDAgger [30]. In the selective strategy, a new neural network, $\pi_{\textsubscript{error}}$, is introduced to estimate the discrepancy between the simulated input and the model’s estimated input. Inputs are provided only at steps with high predicted discrepancy. Mean squared error is used to quantify the discrepancy for feedback, while Kullback–Leibler (KL) divergence is applied for action advice.
@@
11:  Update the $F$ function ($\pi_{\text{grower}}$) and $\pi_{\text{error}}$ (if exists) using $D_{\text{input}}$ for NN iterations
@@ 4.3 Interactive RL Implementation
The implemented interactive RL algorithms are detailed in Algorithm 1. Modifications to the baseline PPO framework for incorporating grower inputs and addressing limited input availability are emphasized in bold. In addition to the introduction of auxiliary neural networks, a new buffer, $D_{\textsubscript{input}}$, is implemented to store data for updating these networks. Furthermore, three new hyperparameters are defined: the weight factor $\beta$, the number of inputs $n$, and the number of update iterations $N$. The weight factor $\beta$ regulates the influence of grower input, as described in (4), (5), and (6). The parameter $n$ constrains the number of inputs provided during each update interval. The number of update iterations $N$ specifies the number of times the auxiliary networks are updated within each interval.
@@ 5.1 Experiment Settings
[…]
@@ 5.3 Performance of Interactive RL
Table 9 shows the average cumulative reward of interactive RL algorithms in the test environment. In all experiments, the availability of inputs is full, and the input weights (likelihood) $\beta$ are chosen to be the ones that work best in the test. Additionally, baseline PPO using only simulated feedback as the reward function is also trained and tested, with its test reward reduced to only 0.71 Hfl/m2. This is expected and indicates that the simulated inputs are far from perfect. The policy shaping using precise action advice performs best in this environment, with 8.4% improvement. Policy shaping using constraints and control sharing also increase rewards. However, reward shaping reduces the reward, indicating that it is more sensitive to the quality of input.
@@
Figure 5 shows the average setpoints and indoor climate variables for each hour of the day. All interactive RL algorithms, except for reward shaping, show a similar pattern. They maintain lower temperature setpoints and consequently lower indoor temperatures than the baseline PPO. Notably, indoor temperatures and CO2 setpoints are generally higher than the corresponding indoor values, and high indoor CO2 and indoor values are observed at night. This is due to several factors: First, the greenhouse environment has no actuator that can actively reduce temperature or CO2 level. Second, in the later stage of the growing cycle, the plant’s respiration can produce a large amount of CO2. Third, since there are no violation constraints on high CO2 level or temperature at night and relative humidity remains within constraints, the agent is not motivated to use ventilation to lower these variables.
@@ 6 Conclusion
$\phi_{\text{transp,h}}(t)={}$ $p_{4,2}(1-\text{exp}(-p_{1,3}x_{1}(t)))$
$\times\left(\frac{p_{4,3}}{p_{4,4}(x_{3}(t)+p_{4,5})}\text{exp}\left(\frac{p_{4,6}x_{3}(t)}{x_{3}(t)+p_{4,7}}\right)-x_{4}(t)\right),$
$u_{1}(t)={}$ $p_{5,1}(u_{CO_{2}}(t)-y_{2}(t))+p_{5,2}\int_{t-1\,\text{hour}}^{t}(u_{CO_{2}}(\tau)-y_{2}(\tau))\,d\tau,$
$u_{3}(t)={}$ $p_{5,3}(u_{temp}(t)-y_{3}(t))+p_{5,4}\int_{t-1\,\text{hour}}^{t}(u_{temp}(\tau)-y_{3}(\tau))\,d\tau,$
@@
$y_{2}(t)$ $=\frac{10^{6}\cdot p_{2,4}(x_{3}(t)+p_{2,5})}{p_{2,6}p_{2,7}}\cdot x_{2}(t),$ $\quad\text{ppm},$
$y_{3}(t)$ $=x_{3}(t),$ $\quad\text{\textdegree C},$
$y_{4}(t)$ $=\frac{10^{2}\cdot p_{2,4}\big{(}x_{3}(t)+p_{2,5}\big{)}}{11\cdot\text{exp}\Big{(}\frac{p_{4,8}x_{3}(t)}{x_{3}(t)+p_{4,9}}\Big{)}}\cdot x_{4}(t),$ $\quad\text{\%},$
@@
$x_{4}(t)$ indoor humidity (%) $d_{4}(t)$ outdoor humidity (kg/m3)
$u_{1}(t)$ CO2 injection (mg/m2/s) $y_{1}(t)$ dry-weight (kg/m2)